OpenStudy (mylahb):

write an explicit formula for each geometric sequence. Then generate the first five terms. a1=6, r=2 @sissyedgar

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_1=6 }\) \(\large\color{black}{ \displaystyle r=2 }\) \(\large\color{black}{ \displaystyle a_2=6\times 2=12 }\) \(\large\color{black}{ \displaystyle a_3=12 \times 2=24 }\)

OpenStudy (solomonzelman):

and on...

OpenStudy (mylahb):

but what is the explicit formula for this equation @SolomonZelman

OpenStudy (solomonzelman):

well \(\large\color{black}{ \displaystyle a_n=\left(a_1\right)\times({\rm r})^{n-1} }\)

OpenStudy (solomonzelman):

use, \(\large\color{black}{ \displaystyle a_n=\left(a_1\right)\times({\rm r})^{n-1} }\) but in your case, \(\large\color{black}{ \displaystyle a_n=\left(6\right)\times({\rm 2})^{n-1} }\)

OpenStudy (mylahb):

\[\sum_{n=1}^{5} (2n+3)\] @SolomonZelman

OpenStudy (mylahb):

find the sum of each finite series

OpenStudy (solomonzelman):

no, that is a series

OpenStudy (solomonzelman):

that is the sum of all terms

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \left\{6\times 2^{n-1}\color{white}{\Huge|}\right\}_{n=1}^{\infty} }\)